OpenStudy (anonymous):

What is the x-intercept of the line with the equation -2x+1/2y=18 ? (__ , 0)

OpenStudy (anonymous):

is there any other part of the question?

OpenStudy (anonymous):

no that is all, you can use this if you need a visual problem https://www.desmos.com/calculator

OpenStudy (anonymous):

never mind i found the answer -9 :) thank you

OpenStudy (lifeisadangerousgame):

x-intercept is when y = to 0. So first, turn this equation into a regular y = mx + b equation. (IN other words, get 'y' by itself on one side!) -2x+1/2y=18 +2x +2x 1/2y = 2x + 18 divide by 1/2 on both sides y = x + 9 So now. set y to equal 0 so: 0 = x + 9 Now solve for x. 0 = x + 9 - 9 - 9 -9 = x So when 'y' is 0, then x is -9. That means, the x-intercept is -9. The answer is (-9,0) Make sense?

OpenStudy (lifeisadangerousgame):

Haha okay!

OpenStudy (anonymous):

as you were saying that i found out how to do it ha :) thank you anyways i medaled you

OpenStudy (lifeisadangerousgame):

Thanks! :D